Reports 1-1 of 1 Clear search Modify search
PEM (Center)
takaaki.yokozawa - 9:36 Friday 06 February 2026 (36301) Print this report
PEM injection test 260206
I performed the PEM injection test (continuously OMC bellows leg injection for evaluating the noise projection below 400 Hz)

07:07:00 - 07:20:00 silent run

07:20:00 - 08:03:40

Swept sine injection, 20 cnt, 400 - 180 Hz, 1 Hz resolution, 10 s in each frequency
K1:PEM-EXCITATION_SR3_RACK_7_EXC
Locked loss happened

08:30:00 - 08:56:00
K1:PEM-EXCITATION_SR3_RACK_7_EXC
Swept sine injection, 20 cnt, 180 - 50 Hz, 1 Hz resolution, 10 s in each frequency

9:00:00 - 9:35:00
K1:PEM-EXCITATION_SR3_RACK_8_EXC
Swept sine injection, 800 cnt, 50 - 10 Hz, 1 Hz resolution, 50 s in each frequency, 1 s setting time, 1 s ramp down
REF : K1:PEM-SEIS_OMC_GND_Z_OUT_DQ or K1:PEM-ACC_OMC_CHAMBER_OMCLEG_Z_OUT_DQ
Locked loss happened during the injection (maybe 12 or 13 Hz)
Comments to this report:
tatsuki.washimi - 12:36 Friday 06 February 2026 (36303) Print this report

Spectrograms for the  small shaker tests.

Images attached to this comment
tatsuki.washimi - 12:51 Friday 06 February 2026 (36304) Print this report

Spectrograms for the  large shaker tests.

Images attached to this comment
Search Help
×

Warning

×